# Prediction and identification of point defect fingerprints in X-ray photoelectron spectra of TiNx with 1.18 1.37

### v2

@@ Abstract
Our recent work showed that it is possible to detect and even quantify the concentration of Ti vacancies in Ti1-δOxN1-x thin films using X-ray photoelectron spectroscopy (XPS) [29]. This was enabled by guiding the XPS analysis with ab initio calculations that allow to precisely predict core electron binding energies (BE) shifts using the density functional theory together with core-hole approaches [30, 31, 32]. Such a combined approach can help to explain spectral features observed experimentally [33] and to obtain important information about the structure of the materials [34]. XPS has already been used to characterize the surfaces [35, 36, 37, 38] and oxidation [39, 40] of TiN films, since even short exposure to air leads to the formation of oxide and oxynitride overlayer at the surface [41, 42]. As the inelastic electron mean free path in TiO2 is just 1.5 nm at 1000 eV [43], the surface oxide makes it difficult to obtain information about the unoxidized TiN coating itself and thus about the point defects in the film. For example, it was previously shown that Ti-deficient TiNx is exhibiting some shift of N 1s binding energies to lower values [36, 44], however, the specific explanation, whether this is a feature of the film itself or just an effect of the oxide layer, was never clarified. While the straightforward solution to eliminate or greatly reduce the oxygen exposure altogether is to perform the XPS measurements in situ [45, 36], recently, capping approaches were developed to prevent oxidation of the TiN films even when the samples are exposed to atmosphere [46, 47]. It was also shown that when the oxidation is small, angle-resolved XPS can be used to distinguish the signal coming from below the oxidized region [45]. Therefore, it is possible to use XPS to obtain information from the bulk-like region of the TiN thin film.
@@ 2.1 Experiment
TiNx thin films were deposited onto Si (100) substrates at floating potential using reactive high-power pulsed magnetron sputtering (HPPMS) of a 50 mm diameter Ti target (99.99 % purity, EVOCHEM Advanced Materials) with an average power density of 9.4 W/cm2, 2.5% duty cycle and 50 s on-time. The depositions were performed in a pure nitrogen atmosphere at a pressure of 1 Pa and a target-to-substrate distance of 10 cm for 2 hours. The thin film composition was controlled by varying the substrate temperature systematically, from = 100 to 400 ∘C in 100 ∘C increments The base pressure at deposition temperature was for all depositions less than Pa. After growth, the as deposited thin films were cooled to room temperature for two hours, while still under vacuum. Subsequently, and before atmosphere exposure, a thin Al capping layer was deposited for 12 s by direct current magnetron sputtering of a diameter-50 mm Al target at 3.9 W/cm2 power density in Ar atmosphere at 0.5 Pa. The Al capping layer was deposited at room temperature since a pronounced island growth was observed at elevated temperatures. Sputter cleaning behind a closed shutter for at least 2 minutes was performed for both Ti and Al targets before the TiN and Al depositions, respectively. After capping and venting, the samples were immediately transferred to the KRATOS AXIS SUPRA X-ray photoelectron spectrometer to limit the surface contamination by atmosphere exposure, keeping the total air exposure time below 10 minutes.
@@ 2.2 Ab initio modelling
The N 1s and Ti 2p3/2 BEs were calculated as the energy difference between the final state, where a single electron was removed from the specified atom and core level and placed in the valence band, and the initial ground state. Self-consistent DFT calculations were used for both the ground and the final state. The final state always corresponds to the lowest energy state under the constraint of the specific core electron being removed from the core level and placed into the valence band. As a result, the method is able to predict the main XPS peaks. However, spectral features corresponding to additional energy loss events, e.g., the satellite peaks, are not modeled. We note that while the absolute BEs do not correspond to the absolute experimental values by few eV, the relative BE shifts using the same method for Ti1-δOxN1-x were previously found to be in perfect agreement with the experiment [29]. The BE calculations were carried out for every atom in the supercell, leading to 64 core hole calculations per defect model and used the all-electron full-potential Wien2k code with the linearized augmented plane-wave basis set [63]. Atomic muffin-tin radii were set to 1.79 and 1.25 Bohr radius for Ti and N atoms, respectively. The basis size parameter of 7.5 and 444 -point grid should guarantee a convergence of better than 0.05 eV for the calculated BEs (while the convergence of the relative energy shifts should be even better). Perdew–Burke–Ernzerhof (PBE) functional was used for all calculations [64]. For the purpose of visualization and comparison with experimental spectra, the discrete calculated binding energies were broadened by Gaussians with eV and the energies were shifted such that the main N 1s peak corresponding to stoichiometric TiN without defects is at 397.3 eV and main Ti 2p3/2 peak at 455.1 eV [35]. All DFT calculations are publicly available under the Creative Commons license in the NOMAD Archive [65, 66].
@@ 3.1 Film composition and structure
The chemical compositions of all as deposited TiNx films was characterized by ion beam analysis and are summarized in Table 1. All measured samples are Ti-deficient with respect to stoichiometric TiN1.0 and Ti content is decreasing with the decreasing deposition temperature. Figure 1 shows XRD patterns of the deposited films. All films show a highly textured cubic TiN phase, however, with the decreasing deposition temperature, the intensity of the diffraction peaks clearly decreases and larger peak broadening is observed, indicating smaller TiN crystallites. Additionally, the peak shifts to smaller angles at lower deposition temperatures, thus corresponding to a lattice parameter increase for films deposited at lower temperatures. However, due to very significant texturing in the films and small grain sizes, the evaluation of stress-free lattice parameter, e.g., by method, was not possible.
@@ 3.3 XPS measurements
Measured Ti 2p XPS spectra are plotted in Figure 3b). The spectra exhibit a structure consisting of four peaks, where the main peaks around 453 eV and 461 eV correspond to the main Ti 2p3/2 and Ti 2p1/2 peaks of TiN, respectively, and the weaker peaks around 458 and 464 eV are the satellites. No other peaks are observed. Ti 2p BEs exhibit smaller changes with the increasing off-stoichiometry compared to the case of N 1s. There is virtually no difference between the TiN1.18 and TiN1.23 samples. As the off-stoichiometry is increased, the main peaks shift to slightly higher BEs and there is a small increase in spectra intensity between the main and satellite peaks.
Based on the XRD, ion beam analysis and XPS results as well as previously reported DFT calculations [16, 67, 55], the following, preliminary conclusions about the point defects potentially present in the films can be drawn. There are three possible point defect-based explanations of N excess as measured by the ion beam analysis: Ti vacancies, N interstitials, and N antisites. Ti vacancies have the lowest energy of formation of 3.2 eV [16, 67]. N interstitials have energy of formation of 4.6, 4.8 and 5.46 eV for the split (10)-aligned, split (111)-aligned, and tetrahedral interstitials [55], respectively. The least energetically favorable defect is the N antisite with formation energy of 12.6 eV [16]. Therefore, based on thermodynamic considerations, Ti vacancies are the most probable point defects facilitating the observed off-stoichiometry, due to their low energy of formation. The shift of the XRD peaks to smaller angles (indicating increasing lattice parameter) for films deposited at lower temperatures with respect to the films deposited at higher temperatures and also to the powder TiN values, i.e., PDF Card No.: 00-038-1420, may indicate the presence of N interstitials. However, since it was not possible to determine the equilibrium lattice parameter, it is possible that the XRD peak shifts are caused by differences in residual stress states of the films. Therefore, based on the ion beam analysis and XRD data, the presence of N interstitials appear possible but cannot be proven.
@@ 3.4.1 Core electron binding energies
In the following, we will exclusively discuss the BEs of N 1s and Ti 2p3/2 electrons. The point defect visualizations, BE histograms and broadened BEs, as well as the mean BE shifts are summarized in Figure 4 for all the here studied point defects. Only atoms with core electron BE shifts ( eV) are marked in the figures. We note that this chosen BE shift limit is to large extend arbitrary and the experimental resolvability of specific spectral components depends on multiple conditions, most notably the instrumental resolution, temperature, crystalline quality (structure disorder), signal-to-noise ratio and intensity ratios of the components. In the case of the here studied films, the FWHM of the main N 1s spectral component is in the range of 0.65–0.8 eV (as shown in Table 3.5 which will be discussed in more detail later) for the measurements at 10 eV pass energy. As a very crude estimate, two components could be resolved if their BE difference is at least half of this FWHM value. Therefore, not all of the bellow discussed BE shifts can be resolved by the spectrometer employed in the present work and furthermore, not all the discussed point defects are expected to be present in the here experimentally studied TiNx with 1.18 1.37. Specifically, presence of Ti interstitials or N vacancies is unlikely, for reasons discussed below. Nevertheless, the following section is an overview of which point defect induced BE shifts are relevant for the experimental analysis of point defects in TiN TiNx (both and ) by XPS, assuming in some cases a better instrumentation, e.g., a synchrotron-based XPS.
@@
N vacancy (Figure 4, first row): The second nearest N2n neighbors of the vacancy see a slight increase in BEs by around 0.29 eV. This is consistent with previous ELNES calculations, where the influence of N vacancy was also reported to be strongest for the second nearest N neighbors [69]. Other N atoms considered here show BE changes eV and are hence not resolvable with laboratory XPs systems. While the N vacancies introduce a positive shift in BEs of its six first Ti1n neighbors, the shift magnitude is only 0.23 eV, and is therefore challenging to be resolved experimentally. The shift of BEs of the second nearest N vacancy N neighbors is larger with almost 0.3 eV. However, the position overlaps with the very broad N 1s satellite peak observed in experiments [35, 42], which is not possible to be modeled by our DFT approach. This complicates fitting and introduces further potential uncertainties.
@@
N tetrahedral interstitial (Figure 4, second row): The interstitial Ni atom exhibits a reduced BE of 1.18 eV, while a BE reduction of 0.53 eV was predicted for its first N1n neighbors. Both BE shifts of eV and eV are resolvable by laboratory XPS systems. All other N and Ti atoms are predicted to exhibit unresolvable BE shifts.
@@
N split (10)-aligned interstitial (Figure 4, third row): The two split interstitial N atoms show a BE increase of 0.46 eV. The first Na neighbors in the direction of the N2 bond exhibit a reduced BE by 0.31 eV. The first Nb neighbors perpendicular to the N2 bond see a BE increase of 0.19 eV. The rest of N atoms exhibits a BE shift smaller than 0.1 eV. The two Tia atoms directly above and below the N2 bond exhibit increase of BE by 0.16 eV, while the remaining four first Tib neighbors see decrease of BE by 0.3 eV. The BE shift of the interstitial N atom of 0.46 eV is resolvable by laboratory XPS, however, the signal is weak as it originates only from two atoms per defect. Potential characterization would be further complicated by an overlap with the satellite peak, which is also situated at higher BE and, therefore, this type of interstitial could be difficult to be unambiguously resolved by fitting extra spectral components.
@@
N split (111)-aligned interstitial (Figure 4, fourth row): There is no N atom with a BE shift eV and the shift of the Ti first neighbors is only eV. Therefore, this type of N interstitial is expected to only lead to small peak broadening and, thus, is deemed to be undetectable by XPS.
@@
N Frenkel pair (Figure 4, fifth row): The calculated Frankel pair defect consists of the N split (10)-aligned interstitial and a close N vacancy in the direction perpendicular to the N–N bond. The calculated BE shifts are very close to what would be expected from just adding together the effects of the before-calculated N vacancy and N split (10)-aligned interstitial. The two interstitial atoms Ni have their N 1s BE increased by 0.47 eV (as compared to 0.46 eV for the separate interstitial). The first Tia and Tib neighbors of the interstitial exhibit Ti shifts of 0.21 eV and eV, respectively, which is similar to the influence of the interstitial alone. A similar effect is seen for the first Tid neighbors of the vacancy, where a BE decrease of 0.23 eV is predicted (identical to the effect of the vacancy alone). Two Tic atoms, which are both first neighbors of the vacancy and interstitial, exhibit a BE shift of eV. In this case, the shift is the exact sum of the effects calculated for N vacancy eV and N interstitial eV.
@@
Ti tetrahedral interstitial (Figure 4, seventh row): This Ti interstitial induces unresolvable changes in the N 1s BEs of the surrounding N atoms. However, the interstitial Tii atom itself exhibits BE reduction by 1.53 eV and its first 4 Ti1n neighbors see a BE reduction of 0.71 eV.
Ti split (10)-aligned interstitial (Figure 4 eighth row): The only significantly influenced N atoms are the nearest Na neighbors in the (100) plane, while no other calculated N 1s binding energies stand out. The general distribution of the remaining BEs from N atoms is quite broad. This is expected as Ti introduces a much larger overall displacements than the smaller N interstitials. The two Tii interstitial atoms show a decrease of binding energy by 1.80 eV, Tia atoms in the direction of the Ti–Ti interstitial pair bond have BEs reduced by 0.43 eV, and the rest of Ti atoms is unaffected.
Ti interstitial 111 (Figure 4, ninth row): The first 6 N1n neighbors of the Ti interstitial pair have their BEs reduced by 0.36 eV. Other N atoms see no significant change. Two Tii interstitial atoms exhibit a BE decrease of 2.28 eV and the Ti1n first neighbors exhibit a BE decrease of 0.28 eV.
Importantly, the predictions highlight that all Ti interstitials exhibit a distinct signal in the Ti 2p spectra. Although the presence of Ti interstitials is quite unlikely due to having very high formation energy of more than eV [67], so N vacancies are more likely in the case of Ti-rich stoichiometry, there are no other overlapping signals at lower BEs with respect to the Ti 2p TiN signal. Therefore, even very small Ti interstitial concentrations should be detectable by laboratory XPS.
In general, the fitting of Ti 2p peak is challenging due to the overlap between the Ti 2p1/2 and Ti 2p3/2 peaks and the presence of satellite peaks which are much more pronounced than in the N 1s signal. The choice of background fitting function and also the choice of a proper asymmetric line shapes is crucial [42]. Based on the discussion above, it is expected that only the Ti interstitials are readily identifiable from the Ti 2p spectra.
Regarding the analysis of the N 1s spectra, we predict a possibility of detecting point defects, which induce a significant negative BE shift of the N 1s atoms, such as Ti vacancies or N tetrahedral interstitials as there is no overlapping peak at lower BEs (assuming the film can be protected from oxidation [42]). Nevertheless, it seems difficult to distinguish Ti vacancies and N tetrahedral interstitials from the XPS spectra. Selected Ti interstitials also induce negative BE shifts in the N 1s spectra, however, BE predictions show that their presence or lack thereof can be easily determined from analyzing the Ti 2p peak. Point defects inducing positive BE shifts of the N 1s core electrons, like N vacancies and N split (10)-aligned interstitials, are predicted to be difficult to detect due to overlap with the satellite peak located at higher BE with respect to the main TiN N 1s component.

### v3

@@ Abstract
Our recent work showed that it is possible to detect and even quantify the concentration of Ti vacancies in Ti1-δOxN1-x thin films using X-ray photoelectron spectroscopy (XPS) [29]. However, only the effect of Ti vacancies was studied in detail. The impact of other possible point defects—such as interstitials, whose presence was also suggested by the XRD results—on the XPS spectra was not thoroughly investigated. This vacancy quantification work was enabled by guiding the XPS analysis with ab initio calculations that allow to precisely predict core electron binding energies (BE) shifts using the density functional theory together with core-hole approaches [30, 31, 32]. Such a combined approach can help to explain spectral features observed experimentally [33] and to obtain important information about the structure of the materials [34]. XPS has already been used to characterize the surfaces [35, 36, 37, 38] and oxidation [39, 40] of TiN films, since even short exposure to air leads to the formation of oxide and oxynitride overlayer at the surface [41, 42]. As the inelastic electron mean free path in TiO2 is just 1.5 nm at 1000 eV [43], the surface oxide makes it difficult to obtain information about the unoxidized TiN coating itself and thus about the point defects in the film. For example, it was previously shown that Ti-deficient TiNx is exhibiting some shift of N 1s binding energies to lower values [36, 44], however, the specific explanation, whether this is a feature of the film itself or just an effect of the oxide layer, was never clarified. While the straightforward solution to eliminate or greatly reduce the oxygen exposure altogether is to perform the XPS measurements in situ [45, 36], recently, capping approaches were developed to prevent oxidation of the TiN films even when the samples are exposed to atmosphere [46, 47]. It was also shown that when the oxidation is small, angle-resolved XPS can be used to distinguish the signal coming from below the oxidized region [45]. Therefore, it is possible to use XPS to obtain information from the bulk-like region of the TiN thin film.
@@ 2.1 Experiment
TiNx thin films were deposited onto Si (100) substrates at floating potential using reactive high-power pulsed magnetron sputtering (HPPMS) of a 50 mm diameter Ti target (99.99 % purity, EVOCHEM Advanced Materials) with an average power density of 9.4 W/cm2, 2.5% duty cycle and 50 s on-time. The depositions were performed in a pure nitrogen atmosphere in order to produce highly Ti-deficient films at a pressure of 1 Pa and a target-to-substrate distance of 10 cm for 2 hours. The reason for using pure nitrogen as opposed to more common Ar+N2 mixture was to get as N-rich film as possible and therefore large Ti vacancy concentrations. Some of the films with smaller off-stoichiometry could have been deposited also from the mixed Ar+N2 atmosphere, however, pure nitrogen was used for all samples for consistency. The thin film composition was controlled by varying the substrate temperature systematically, from = 100 to 400 ∘C in 100 ∘C increments The base pressure at deposition temperature was for all depositions less than Pa. After growth, the as deposited thin films were cooled to room temperature for two hours, while still under vacuum. Subsequently, and before atmosphere exposure, a thin Al capping layer was deposited for 12 s by direct current magnetron sputtering of a diameter-50 mm Al target at 3.9 W/cm2 power density in Ar atmosphere at 0.5 Pa. The Al capping layer was deposited at room temperature since a pronounced island growth was observed at elevated temperatures. Sputter cleaning behind a closed shutter for at least 2 minutes was performed for both Ti and Al targets before the TiN and Al depositions, respectively. Al capping was chosen based on the work by Greczynski et al. [46], since Al forms a dense oxide scale with only 4 nm thickness even after prolonged exposure to air [48] and is a light element, so XPS signal attenuation is low. After capping and venting, the samples were immediately transferred to the KRATOS AXIS SUPRA X-ray photoelectron spectrometer to limit the surface contamination by atmosphere exposure, keeping the total air exposure time below 10 minutes.
@@
Film thicknesses were measured using a FEI Helios Nanolab 660 dual-beam microscope. The region of interest was marked with a Pt protection layer and a trench was milled, employing focused ion beam techniques with Ga ions at 30 kV. The film thicknesses were determined at an angle of 52 degrees with tilt correction.
@@ 2.2 Ab initio modelling
The N 1s and Ti 2p3/2 BEs were calculated as the energy difference between the final state, where a single electron was removed from the specified atom and core level and placed in the valence band, and the initial ground state, . Self-consistent DFT calculations were used for both the ground and the final state. The final state always corresponds to the lowest energy state under the constraint of the specific core electron being removed from the core level and placed into the valence band. As a result, the method is able to predict the main XPS peaks. However, spectral features corresponding to additional energy loss events, e.g., the satellite peaks, are not modeled. We note that while the absolute BEs do not correspond to the absolute experimental values by few eV, the predicted N 1s BEs of the TiN main component are at eV instead of 397.3 eV [35], the relative BE shifts using the same method for Ti1-δOxN1-x were previously found to be in perfect agreement with the experiment [29]. The differences in absolute binding energies are to some extent caused by the Coulombic interaction between the core hole and its periodic images. However, the impact of this effect on relative shifts is assumed to be small due to the good electrical conductivity and efficient screening in TiN. The BE calculations were carried out for every atom in the supercell, leading to 64 core hole calculations per defect model and used the all-electron full-potential Wien2k code with the linearized augmented plane-wave basis set [64]. The predicted binding energy shift for atoms in the vicinity of a point defect was calculated as the difference between their BEs and the average BE of all other atoms of the same type in the structure. As an example, the N 1s BEs of the six equivalent N neighbors around a Ti vacancy were compared to the mean N 1s BE value of the remaining 26 N atoms in the supercell. The reference BEs, labeled TiN, are shown in red in the histograms of Figure 4. Atomic muffin-tin radii were set to 1.79 and 1.25 Bohr radius for Ti and N atoms, respectively. The basis size parameter of 7.5 and 444 -point grid should guarantee a convergence of better than 0.05 eV for the calculated BEs (while the convergence of the relative energy shifts should be even better). Perdew–Burke–Ernzerhof (PBE) functional was used for both VASP and Wien2k calculations [65]. For the purpose of visualization and comparison with experimental spectra, the discrete calculated binding energies were broadened by Gaussians with eV. We note that the switch between two different DFT codes was done for efficiency purposes. VASP offers fast structure relaxation and has a NEB implementation. Wien2k is a slower all-electron code, however, it offers explicit modeling of the core states which benefits the core-hole calculations, e.g., fully relativistic and self-consistent treatment of all core electrons by default. All DFT calculations are publicly available under the Creative Commons license in the NOMAD Archive [66, 67].
@@ 3.1 Film composition and structure
The chemical compositions of all as deposited TiNx films was characterized by ion beam analysis and are summarized in Table 1. All measured samples are Ti-deficient with respect to stoichiometric TiN1.0 and Ti content is decreasing with the decreasing deposition temperature. Figure 1 shows XRD patterns of the deposited films. All films show a highly textured cubic TiN phase, however, with the decreasing deposition temperature, the intensity of the diffraction peaks clearly decreases and larger peak broadening is observed, indicating smaller TiN crystallites. Additionally, the peak shifts to smaller angles at lower deposition temperatures, thus corresponding to a lattice parameter increase for films deposited at lower temperatures. However, due to very significant texturing in the films and small grain sizes, the evaluation of stress-free lattice parameter, e.g., by method, was not possible. The film thicknesses varied between 120 and 90 nm (as determined from film cross section measurements shown in Supplementary Information, Figure SI3) corresponding to deposition rates 1.0–0.7 nm/min.
@@ 3.3 XPS measurements
Measured Ti 2p XPS spectra are plotted in Figure 3b). The spectra exhibit four visible structures, where the main features around 455 eV and 461 eV are attributed to the main Ti 2p3/2 and Ti 2p1/2 peaks of TiN, respectively, and the weaker features around 458 and 464 eV are satellites [35]. No other structures are observed. Ti 2p BEs exhibit smaller changes with the increasing off-stoichiometry compared to the case of N 1s. There is virtually no difference between the TiN1.18 and TiN1.23 samples. As the off-stoichiometry is increased, the main peaks shift to slightly higher BEs and there is a small increase in spectra intensity between the main and satellite peaks. Measured XPS overview spectra as well O 1s, Al 2p, C 1s and valence band regions spectra are show in the Supplementary Information, Figures SI1 and SI2. We note that the observed oxygen and carbon signals, of varying magnitude from sample to sample, are in excellent agreement with reactions taking place on the Al capping layer: The O 1s binding energies in the range of 532.2–532.4 eV is consistent with the formation of Al–O bonds at 532.0–532.4 eV [46])while the C–C/C–H C 1s peak at 286.1–286.3 eV conforms to expected BE of adventitious carbon on air-exposed Al at 286.1-–286.2 eV [68]), confirming functionality of the capping layer. Furthermore, there is no evidence for the incorporation of C into the TiN as Ti–C components in the C 1s spectra 282 eV [69, 70, 71] are absent, see Figure SI2b. Therefore, the random variations in the C 1s and O 1s intensities are ascribed to the differences in the ambient conditions and air exposure time during venting and sample transport procedures and not to compositional changes in the TiN layer.
Based on the XRD, ion beam analysis and XPS results as well as previously reported DFT calculations [16, 72, 56], the following, preliminary conclusions about the point defects potentially present in the films can be drawn. There are three possible point defect-based explanations of N excess as measured by the ion beam analysis: Ti vacancies, N interstitials, and N antisites. Ti vacancies have the lowest energy of formation of 3.2 eV [16, 72]. N interstitials have energy of formation of 4.6, 4.8 and 5.46 eV for the split (10)-aligned, split (111)-aligned, and tetrahedral interstitials [56], respectively. The least energetically favorable defect is the N antisite with formation energy of 12.6 eV [16]. Therefore, based on thermodynamic considerations, Ti vacancies are the most probable point defects facilitating the observed off-stoichiometry, due to their low energy of formation. The shift of the XRD peaks to smaller angles (indicating increasing lattice parameter) for films deposited at lower temperatures with respect to the films deposited at higher temperatures and also to the powder TiN values, i.e., PDF Card No.: 00-038-1420, may indicate the presence of N interstitials. However, since it was not possible to determine the equilibrium lattice parameter, it is possible that the XRD peak shifts are caused by differences in residual stress states of the films, i.e., larger residual stress is to be expected in the films deposited at higher temperatures. Therefore, based on the ion beam analysis and XRD data, the presence of N interstitials appear possible but cannot be definitely proven.
@@ 3.4.1 Core electron binding energies
In the following, we will exclusively discuss the BEs of N 1s electrons. The point defect visualizations, BE histograms and broadened BEs, as well as the mean BE shifts are summarized in Figure 4 for all the here studied point defects. Only atoms with core electron BE shifts ( eV) are marked in the figures. We note that this chosen BE shift limit is to large extend arbitrary and the experimental resolvability of specific spectral components depends on multiple conditions, most notably the instrumental resolution, temperature, crystalline quality (structure disorder), signal-to-noise ratio and intensity ratios of the components. In the case of the here studied films, the FWHM of the main N 1s spectral component is in the range of 0.65–0.8 eV (as shown in Table 3 which will be discussed in more detail later) for the measurements at 10 eV pass energy. As a very crude estimate, two components could be resolved if their BE difference is at least half of this FWHM value. Therefore, not all of the bellow discussed BE shifts can be resolved by the spectrometer employed in the present work and furthermore, not all the discussed point defects are expected to be present in the here experimentally studied TiNx with 1.18 1.37. Specifically, presence of Ti interstitials or N vare presented in supplementary informationacancies is unlikely, for reasons discussed below. Nevertheless, the following section is an overview of which point defect-induced BE shifts are relevant for the experimental analysis of point defects in TiNx (both and ) by XPS, assuming in some cases a better instrumentation, e.g., a synchrotron-based XPS. While our experimental analysis utilizes only predictions of the N 1s BEs, we have also calculated the Ti 2p3/2 BEs. We propose their potential usefulness for detecting Ti interstitials; however, a rigorous experimental validation is beyond the scope of the present work. Interested reader can find the calculated BEs in the supplementary information as Figure SI5, together with brief discussion.
@@
N vacancy (Figure 4, first row): The second nearest N2n neighbors of the vacancy see a slight increase in BEs by around 0.29 eV. This is consistent with previous ELNES calculations, where the influence of N vacancy was also reported to be strongest for the second nearest N neighbors [74]. Other N atoms considered here show BE changes eV and are hence not resolvable with laboratory XPS systems. The shift of BEs of the second nearest N vacancy N neighbors is larger with almost 0.3 eV. However, the position overlaps with the very broad N 1s satellite peak observed in experiments [35, 42], which is not possible to be modeled by our DFT approach. This complicates fitting and introduces further potential uncertainties.
@@
N tetrahedral interstitial (Figure 4, second row): The interstitial Ni atom exhibits a reduced BE of 1.18 eV, while a BE reduction of 0.53 eV was predicted for its first N1n neighbors. Both BE shifts of eV and eV are resolvable by laboratory XPS systems. All other N atoms are predicted to exhibit unresolvable BE shifts.
@@
N split (10)-aligned interstitial (Figure 4, third row): The two split interstitial N atoms show a BE increase of 0.46 eV. The first Na neighbors in the direction of the N2 bond exhibit a reduced BE by 0.31 eV. The first Nb neighbors perpendicular to the N2 bond see a BE increase of 0.19 eV. The rest of N atoms exhibits a BE shift smaller than 0.1 eV. The BE shift of the interstitial N atom of 0.46 eV is resolvable by laboratory XPS, however, the signal is weak as it originates only from two atoms per defect. Potential characterization would be further complicated by an overlap with the satellite peak, which is also situated at higher BE and, therefore, this type of interstitial could be difficult to be unambiguously resolved by fitting extra spectral components.
@@
N split (111)-aligned interstitial (Figure 4, fourth row): There is no N atom with a BE shift eV, therefore, this type of N interstitial is expected to only lead to small peak broadening and, thus, is deemed to be undetectable by XPS.
@@
N Frenkel pair (Figure 4, fifth row): The calculated Frankel pair defect consists of the N split (10)-aligned interstitial and a close N vacancy in the direction perpendicular to the N–N bond. The calculated BE shifts are very close to what would be expected from just adding together the effects of the before-calculated N vacancy and N split (10)-aligned interstitial. The two interstitial atoms Ni have their N 1s BE increased by 0.47 eV (as compared to 0.46 eV for the separate interstitial).
@@
Ti tetrahedral interstitial (Figure 4, seventh row): This Ti interstitial induces unresolvable changes in the N 1s BEs of the surrounding N atoms.
Ti split (10)-aligned interstitial (Figure 4 eighth row): The only significantly influenced N atoms are the nearest Na neighbors in the (100) plane, while no other calculated N 1s binding energies stand out. The general distribution of the remaining BEs from N atoms is quite broad. This is expected as Ti introduces a much larger overall displacements than the smaller N interstitials.
Ti interstitial 111 (Figure 4, ninth row): The first 6 N1n neighbors of the Ti interstitial pair have their BEs reduced by 0.36 eV. Other N atoms see no significant change.
Therefore, we predict a possibility of detecting point defects which induce a significant negative BE shift of the N 1s atoms, such as Ti vacancies or N tetrahedral interstitials as there is no overlapping peak at lower BEs (assuming the film can be protected from oxidation [42]). Nevertheless, it seems difficult to distinguish Ti vacancies and N tetrahedral interstitials from the XPS spectra. Selected Ti interstitials also induce negative BE shifts in the N 1s spectra, however, they are unlikely based on their high formation energy of more than eV [72], compared to 2.4–2.5 eV for N vacancy [56, 72]. Point defects inducing positive BE shifts of the N 1s core electrons, like N vacancies and N split (10)-aligned interstitials, are predicted to be difficult to detect due to overlap with the satellite peak located at higher BE with respect to the main TiN N 1s component. The effect of possible stress in the films is not taken into account in the first principles BE simulations, they are all done with fully relaxed structures. We have however performed additional calculations, which show that the effect is negligible. Specifically, for the structure with single Ti vacancy with a 0.02 A lattice parameter difference, the change in both absolute BEs and also the BE shifts is less than 0.01 eV, which we deem undetectable by table-top XPS.